OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 4x - 3y = -12 and passes through the point (-3, 4). y = four thirdsx + 8 y = four thirdsx + 3 y = -three fourthsx + 8 y = -three fourthsx + 3

OpenStudy (chosenmatt):

@wolfmaster I think its D

OpenStudy (anonymous):

Um why do you think its D?

OpenStudy (igreen):

@chosenmatt Your answer is wrong.

OpenStudy (chosenmatt):

oh sorry

OpenStudy (igreen):

Parallel lines have the same slope..and the slope of 4x - 3y = -12 is 4/3. So A and B are your only 2 choices..now let's plug in (-3, 4) into each of these equations: y = 4/3x + 8 and y = 4/3x + 3 ----------------- y = 4/3x + 8 4 = 4/3(-3) + 8 Multiply: 4 = -4 + 8 Can you add -4 + 8? @yoyogators

OpenStudy (natsent06):

@iGreen matt is just repeating what he is saying on every question

OpenStudy (igreen):

@yoyogators Are you there? Can you add -4 + 8?